### v1

@@ 3.1 Dataset Collection and Automatic Bone labeling
The key principle of our optimization method is to add small perturbations ($\boldsymbol{\theta}_{\epsilon}$ and $\mathbf{d}_{\epsilon}$) to the ultrasound image poses in order to maximize the intensity of pixels at the intersection between the ultrasound image and bone CT model while minimizing deviations from the recorded tracking data. We define the perturbed point cloud as $\mathcal{U}^{t}(\boldsymbol{\theta}_{\epsilon},\mathbf{d}_{\epsilon})=\mathsf{T}^{t+\Delta t}_{CT\leftarrow US}(\boldsymbol{\theta}+\boldsymbol{\theta}_{\epsilon},\mathbf{d}+\mathbf{d}_{\epsilon})\cdot\mathsf{T}_{US\leftarrow I}\cdot\mathsf{T}_{S}\cdot\mathcal{P}^{t}$ and the perturbed intersection as $\mathcal{I}^{t}(\boldsymbol{\theta}_{\epsilon},\mathbf{d}_{\epsilon})=\mathcal{C}\cap\mathcal{U}^{t}(\boldsymbol{\theta}_{\epsilon},\mathbf{d}_{\epsilon})$. The proposed solution is written as:
@@
$\operatorname*{arg\,max}_{\boldsymbol{\theta}_{\epsilon},\mathbf{d}_{\epsilon}\in[-1,1]^{3}}{\frac{\sum_{\mathbf{u}\in\mathcal{I}^{t}(\boldsymbol{\theta}_{\epsilon},\mathbf{d}_{\epsilon})}{f(\mathbf{u})}}{|\mathcal{I}^{t}(\boldsymbol{\theta}_{\epsilon},\mathbf{d}_{\epsilon})|}-\lambda\frac{\sum_{i}{||\mathsf{T}(\boldsymbol{\theta},\mathbf{d})\cdot\mathbf{F}_{i}-\mathsf{T}(\boldsymbol{\theta}+\boldsymbol{\theta}_{\epsilon},\mathbf{d}+\mathbf{d}_{\epsilon})\cdot\mathbf{F}_{i}||_{2}}}{N}}$ (3)
@@
$\mathbf{F}_{i}$ denotes the homogeneous coordinate vector of the $i$-th tracking fiducial of the ultrasound probe marker, with $N$ fiducials in total. The first term of equation (3) represents the mean intensity values of points in $\mathcal{I}^{t}(\boldsymbol{\theta}_{\epsilon},\mathbf{d}_{\epsilon})$ under the corrected image pose. The second term represents the mean deviation of the tracking fiducials induced by the pose correction in terms of Euclidean distance. By leveraging the bounded set and regularization term, our approach ensures the solution remains close to the initial tracking data. The weighting parameter $\lambda$ is empirically determined from the range [0.01, 0.1, 1, 10, 100,1000]. The optimization problem is solved using the differential evolution method [24]. Each $\lambda$ value was tested on randomly sampled frames, and the results were evaluated quantitatively to select the optimal $\lambda$. To identify failed optimizations, a validation step is performed. Specifically, ultrasound images with a mean fiducial location correction greater than 1 mm after optimization are discarded, as the tracking device provides sub-millimetric accuracy. After the validation step, approximately 18.7% frames are discarded, with the mean fiducial location correction of the remaining frames being 0.481 mm. An example of a full bone label refined using the proposed optimization is shown in Figure 2.
To account the bone shadow effect, bone regions within the shadow (blue region in Figure 2) are discarded. In addition, we project the normal vectors from the bone CT model onto the ultrasound images and compute the incidence angles $\alpha$ between the projected normal vectors and the vertical scan lines, as shown in Figure 2. According to ultrasound imaging physics, the intensity of bone pixels decreases as $\alpha$ increases, with $\alpha>90\degree$ being a critical boundary. In this study, we keep bone pixels with $\alpha\leq 85\degree$ to avoid boundary cases. An example of final bone label is shown in Figure 2. More examples are provided in the supplementary material.
@@
To clinically evaluate the quality of bone annotations following our optimization pipeline, an expert physician specialized on orthopedic sonography assessed the quality of both the initial and optimized bone annotations. For comparison with state-of-the-art, we also employed the traditional ”ultrasound to CT registration” algorithm available in the ImFusion Suite [^6] https://www.imfusion.com/ (ImFusion, Munich, Germany) to compute a global transformation matrix for registering the ultrasound data to the CT volume. In our experiment, one frame was randomly sampled from each ultrasound sweep, resulting in a total of 119 frames. Bone annotations were performed using three alignment methods: the initial alignment, the optimized alignment derived from our method, and the bone alignment obtained after the CT-US registration step. The resulting bone alignment outputs from these three methods were presented to an expert physician specialized on orthopedic sonography in random order. The expert evaluated the quality of each bone alignment using the following criteria:
@@ 3.2 Bone Segmentation
Segmentation Mode. A U-net with ResNet encoders is trained on UltraBones100k to automatically segment bones in ultrasound images. The segmentation loss function being used is a combination of Dice Loss and Binary Cross Entropy Loss (implementation details are provided in the supplementary material). We split the dataset into 10 training specimens, 1 validation specimen, and 3 test specimens (around 25k ultrasound images). The training and validation sets are further split using cross-validation. For the test set, 700 randomly sampled ultrasound images were manually labeled by an expert surgeon, who identified the hyperechoic bone surface layer and its accompanying acoustic shadow using anatomical knowledge and real-time feedback. (detailed in the supplementary material)
@@
Evaluation Metrics. In previous studies [14, 15], over 18 metrics were used for evaluating bone segmentation in ultrasound images, including intersection-based (e.g., Dice score) and distance-based metrics (e.g., one-sided CD), which complicates fair benchmarking of models. Bone labels in ultrasound are thin, even small shifts can drastically reduce intersection-based scores. While CD is easy to interpret, it struggles with false positives. To address this, we propose using distance-thresholded accuracy, completeness, and F1-score, which balance intersection-based and distance-based evaluations. Accuracy represents the fraction of correct predictions out of all predictions at a certain distance threshold, while completeness measures the fraction of correct predictions relative to the ground truth. Explicit formulations of these metrics are provided in the supplementary material. Particularly, our evaluation is based on the CT-derived labels, allowing to compare the performance between automatic segmentation and manual labeling.
@@
To particularly assess the performance of our method in segmenting lower-intensity bone regions, we quantitatively divide the labels into high-intensity and low-intensity regions. The intensity values of all bone pixels in the collected ultrasound images are recorded, and the Otsu method is applied to determine an intensity threshold separating the bone pixels into two classes: high-intensity and low-intensity. The threshold is chosen to minimize intraclass variance and maximize interclass variance. Exemplar results are shown in Figure 3.
4 Results and Discussion
@@
With our methodology, we have collected a dataset of 100k annotated ultrasound images with bone labels from 14 human cadaveric lower legs. The specimens have different physical characteristics, with heights ranging from 155 to 185 cm, weights from 59 to 99 kg, and BMI values between 19 and 30. The age range spans from 41 to 76 years, and the dataset includes both male and female specimens. Furthermore, arthritis status is documented, with some specimens having a history of falls or related conditions.
For the bone annotation rating experiment, the mean ratings are 1.45 for the initial alignment, 0.87 for CT-US registration, and 2.39 for our optimized alignment (box plot available in the supplementary material). A Wilcoxon signed-rank test with Bonferroni correction demonstrated a significant improvement in bone labeling quality following our optimization pipeline ($p<0.001$). Unlike the traditional CT-US registration method, which calculates a single global transformation matrix for the entire ultrasound sweep to maximize overlapping voxel intensities, our approach computes a transformation matrix for each individual ultrasound frame. By leveraging the initial bone alignment as a prior, our method achieves significantly improved bone annotations.
Qualitative results between the manual bone labels and the model-predicted labels are shown in Figure 3 (additional results are provided in the supplementary material). The results indicate that the model trained on UltraBones100k outperforms the surgeon in segmenting bones in ultrasound images. Manual labels mostly include only high-intensity bone regions. In contrast, predicted bone labels include both high-intensity and low-intensity regions.
@@
Quantitative results, including accuracy, completeness, and F1 score, are presented in Figure 4, demonstrating that the trained model consistently outperforms surgeon’s manual labeling, regardless of the distance threshold used. On average, at a distance threshold of 0.5 mm, the trained model achieves an accuracy of 0.900, completeness of 0.895, and an F1 score of 0.893 across all bone regions. Particularly, for low-intensity regions, the accuracy of the trained model (0.789) is 27.4% higher than that of manual labeling (0.619), the completeness (0.801) is 320% higher compared to manual labeling (0.190), and the F1 score (0.778) is 197% higher than the manual labeling (0.261). These results in combination with the intensity distribution of labels illustrated in Figure 3(b) confirm that the expert surgeon faces difficulty in correctly labeling bone contours in low-intensity regions. These findings underscore the superiority of our approach over the manual labeling.
Previous work [20] aimed to align a tracked CT bone model with tracked ultrasound images. However, it was found that a mismatch of 1–2 mm was too large to allow for accurate automatic bone labeling in ultrasound images, necessitating manual bone labeling. In contrast, our optimization pipeline enhances CT-ultrasound alignment, facilitating automatic accurate bone label generation.
Generalizability The advantage of our method is that CT images are required only during pretraining. As demonstrated, this pretraining can even be conducted on ex-vivo anatomy, ensuring that no patients are exposed to ionizing radiation with our approach. Therefore it will not affect the current clinical workflow. Examples of ultrasound images depicting the pelvis and femur from a new specimen, along with the corresponding predicted bone labels by the model pretrained on UltraBones100k, are presented in the supplementary material, which demonstrate the generalizability of the pretrained model. The generalizability arises from the relatively consistent appearance of bones in ultrasound images. Bones, surrounded by soft tissues, exhibit highest acoustic impedance, appearing as bright curves or lines. Unlike other imaging methods, 2D ultrasound captures only partial bone segments, often unaffected by overall anatomical shape. Shared structural features across bones, such as in the arms and legs, further enhance model generalization. However, depending on the required level of accuracy, additional data collection may be necessary for specific anatomies and pathologies. In some cases, this may include collecting of in-vivo data. Conducting data collection studies with a limited number of patients is a common and justifiable approach in such scenarios.

### v2

@@ 3.1 Dataset Collection and Automatic Bone labeling
The key principle of our optimization method is to add small perturbations ($\bm{\theta}_{\epsilon}$ and $\mathbf{d}_{\epsilon}$) to the ultrasound image poses in order to maximize the intensity of pixels at the intersection between the ultrasound image and bone CT model while minimizing deviations from the recorded tracking data. We define the perturbed point cloud as $\mathcal{U}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})=\mathsf{T}^{t+\Delta t}_{\text{CT}\leftarrow\text{US}}(\bm{\theta}+\bm{\theta}_{\epsilon},\mathbf{d}+\mathbf{d}_{\epsilon})\cdot\mathsf{T}_{\text{US}\leftarrow\text{IP}}\cdot\mathsf{T}_{\text{S}}\cdot\mathcal{P}^{t}$ and the perturbed intersection as $\mathcal{I}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})=\mathcal{C}\cap\mathcal{U}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})$. The proposed solution is written as:
@@
$\operatorname*{arg\,max}_{\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon}\in[-1,1]^{3}}{\frac{\sum_{\mathbf{u}\in\mathcal{I}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})}{f(\mathbf{u})}}{|\mathcal{I}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})|}-\lambda\frac{\sum_{i}{||\mathsf{T}(\bm{\theta},\mathbf{d})\cdot\mathbf{F}_{i}-\mathsf{T}(\bm{\theta}+\bm{\theta}_{\epsilon},\mathbf{d}+\mathbf{d}_{\epsilon})\cdot\mathbf{F}_{i}||_{2}}}{N}}$ (2)
@@
$\mathbf{F}_{i}$ denotes the homogeneous coordinate vector of the $i$-th tracking fiducial of the ultrasound probe marker, with $N$ fiducials in total. The first term of equation (2) represents the mean intensity values of points in $\mathcal{I}^{t}(\bm{\theta}_{\epsilon},\mathbf{d}_{\epsilon})$ under the corrected image pose. The second term represents the mean deviation of the tracking fiducials induced by the pose correction in terms of Euclidean distance. By leveraging the bounded set and regularization term, our approach ensures the solution remains close to the initial tracking data. The weighting parameter $\lambda$ is empirically determined from the range [0.01, 0.1, 1, 10, 100,1000]. The optimization problem is solved using the differential evolution method [31]. Each $\lambda$ value was tested on randomly sampled frames, and the results were evaluated quantitatively to select the optimal $\lambda$. To identify failed optimizations, a validation step is performed. Specifically, ultrasound images with a mean fiducial location correction greater than 1 mm after optimization are discarded, as the tracking device provides sub-millimetric accuracy. After the validation step, approximately 18.7% frames are discarded, with the mean fiducial location correction of the remaining frames being 0.481 mm. An example of a full bone label refined using the proposed optimization is shown in Figure 3. To account the bone shadow effect, bone regions within the shadow (blue region in Figure 3) are discarded. In addition, we project the normal vectors from the bone CT model onto the ultrasound images and compute the incidence angles $\alpha$ between the projected normal vectors and the vertical scan lines, as shown in Figure 3. According to ultrasound imaging physics, the intensity of bone pixels decreases as $\alpha$ increases, with $\alpha>90\degree$ being a critical boundary. In this study, we keep bone pixels with $\alpha\leq 85\degree$ to avoid boundary cases. The final bone label is shown in Figure 3. More examples of bone labels before and after the optimization are shown in Figure 4.
Figure 4: Comparison results of bone annotation before and after optimization. Red regions show the initial alignment, green regions show the refined alignment, yellow regions are the overlap.
@@
Clinical Evaluation. To clinically evaluate the quality of bone annotations following our optimization pipeline, an expert physician specialized on orthopedic sonography assessed the quality of both the initial and optimized bone annotations. For comparison with state-of-the-art, we also employed the traditional ”ultrasound to CT registration” algorithm available in the ImFusion Suite [^4] https://www.imfusion.com/ (ImFusion, Munich, Germany) to compute a global transformation matrix for registering the ultrasound data to the CT volume (the default parameter setting yielded the best results in our experiments), rather than optimizing the alignment for each frame as our method does. In our experiment, one frame was randomly sampled from each ultrasound sweep, resulting in a total of 119 frames. Bone annotations were performed using three alignment methods: the initial alignment, the optimized alignment derived from our method, and the bone alignment obtained after the CT-US registration step. The resulting bone alignment outputs from these three methods were presented to an expert physician specialized on orthopedic sonography in random order. The expert evaluated the quality of each bone alignment using the following criteria:
  • 1.
@@
A screenshot of the rating software interface is shown in Figure 5. A scale indicating 1, 2, and 3 mm is displayed in the top-left corner of each image.
Figure 5: A screenshot of the bone alignment rating software.
@@ 3.2 Bone Segmentation
Segmentation Mode. In our study, 14 cadavers were scanned, resulting a dataset of over 100k ultrasound images with bone labels. We split the dataset into 10 training specimens, 1 validation specimen, and 3 test specimens (around 25k ultrasound images). The training and validation sets are further split using cross-validation. A U-net with ResNet encoders is trained on UltraBones100k to automatically segment bones in ultrasound images[32, 33]. The loss function is a weighted combination of Dice Loss (weight=1) and Binary Cross Entropy Loss (weight=1). We trained the model using the Adam optimizer with $\beta=(0.9,0.999)$, starting with an initial learning rate of 1e-5, which decayed by 0.9 every 10 epochs. The training process ran for 100 epochs on an NVIDIA V100 GPU, with a batch size of 32. The code and pretrained model will be publicly available.
Manual annotation. To compare the automatic bone labeling by the trained model with manual labeling, 700 ultrasound images were randomly sampled from the test set and manually labeled by an expert surgeon. The surgeon identified the hyperechoic bone surface layer and its corresponding acoustic shadow based on the following knowledge about ultrasound imaging and anatomical structures. Bone surfaces in ultrasound images are identified as a hyperechoic (bright) layer resulting from the significant reflection of ultrasound waves at the bone-tissue interface. This bright layer is typically accompanied by a characteristic acoustic shadow extending posteriorly, as the underlying bone attenuates or reflects the majority of the ultrasound signal. The anatomical region being scanned significantly influences the clarity and confidence of bone surface identification. Larger bones (e.g., $>$1 cm in diameter) with smooth, superficial surfaces are more readily identified due to their distinct echogenicity and reduced interference from surrounding tissue artifacts or acoustic noise. In contrast, smaller or deeper bones, or those with irregular surfaces, may present more ambiguous features. Experienced operators rely on a combination of real-time image feedback and their knowledge of anatomical landmarks to delineate tissue structures, including bone surfaces. This process underscores the operator-dependent nature of ultrasound-based diagnostics in clinical practice. For labeling purposes, the ImFusion Label software [^5] https://www.imfusion.com/ (ImFusion, Munich, Germany) was employed to annotate the bone surface layer on a pixel-wise basis within the acquired images. This software allows precise manual delineation of the bone surface, facilitating the creation of ground-truth labels for subsequent analysis.”
@@
Evaluation Metrics. In previous studies [21, 22], over 18 different metrics were used for evaluating bone segmentation in ultrasound images, including intersection-based (e.g., Dice score) and distance-based metrics (e.g., one-sided CD), which complicates fair benchmarking of models. Bone labels in ultrasound are thin, even small shifts can drastically reduce intersection-based scores. While CD is easy to interpret, it struggles with false positives. To address this, we propose using distance-thresholded accuracy, completeness, and F1-score, which balance intersection-based and distance-based evaluations. Accuracy represents the fraction of correct predictions out of all predictions at a certain distance threshold, while completeness measures the fraction of correct predictions relative to the ground truth. Particularly, our evaluation is based on the CT-derived labels, allowing to compare the performance between automatic segmentation and manual labeling. Specifically, for a pixel in the predicted bone label $\hat{\mathbf{x}}\in\mathcal{P}$, its distance to the ground truth $\mathcal{G}$ is defined as follows:
$d_{\hat{\mathbf{x}}\rightarrow\mathcal{G}}=\underset{\mathbf{x}\in\mathcal{G}}{\operatorname{min}}\|\hat{\mathbf{x}}-\mathbf{x}\|,$ (3)
The accuracy, completeness, and F1 score with a distance threshold $\sigma$ are defined as follows:
$Acc(\sigma)=\frac{1}{|\mathcal{P}|}\sum_{\hat{\mathbf{x}}\in\mathcal{P}}[d_{\hat{\mathbf{x}}\rightarrow\mathcal{G}}<\sigma]$ (4)
$Com(\sigma)=\frac{1}{|\mathcal{G}|}\sum_{\mathbf{x}\in\mathcal{G}}[d_{\mathbf{x}\rightarrow\mathcal{P}}<\sigma]$ (5)
$F(\sigma)=\frac{2Acc(\sigma)\cdot Com(\sigma)}{Acc(\sigma)+Com(\sigma)},$ (6)
where [.][.][ . ] is the Iverson bracket.
@@
To particularly assess the performance of our method in segmenting lower-intensity bone regions, we quantitatively divide the labels into high-intensity and low-intensity regions. The intensity values of all bone pixels in the collected ultrasound images are recorded, and the Otsu method is applied to determine an intensity threshold separating the bone pixels into two classes: high-intensity and low-intensity [34]. The threshold is chosen to minimize intraclass variance and maximize interclass variance. Exemplary results are shown in Figure 6.
4 Results
((a))
((b))
@@ 4 Results
Dataset Variability. With our methodology, we have collected a dataset of 100k annotated ultrasound images with bone labels from 14 human cadaveric lower legs from different donors. The donors cover different physical characteristics, with heights ranging from 155 to 185 cm, weights from 59 to 99 kg, and BMI values between 19 and 30. The age range spans from 41 to 76 years, and the dataset includes both male and female specimens. Furthermore, arthritis status is documented, with some specimens having a history of falls or related conditions.
Figure 7: Results of the rating experiment.
Clinical Evaluation. The results of the bone annotation rating experiment is shown in Figure 7. The mean ratings are 1.45 for the initial alignment, 0.87 for CT-US registration, and 2.39 for our optimized alignment. A Wilcoxon signed-rank test with Bonferroni correction demonstrated a significant improvement in bone labeling quality following our optimization pipeline ($p<0.001$).
Automatic vs Manual Labeling. Qualitative results between the manual bone labels and the model-predicted labels are shown in Figure 6. Quantitative results, including accuracy, completeness, and F1 score, are presented in Figure 8. On average, at a distance threshold of 0.5 mm, the trained model achieves an accuracy of 0.900, completeness of 0.895, and an F1 score of 0.893 across all bone regions. Particularly, for low-intensity regions, the accuracy of the trained model (0.789) is 27.4% higher than that of manual labeling (0.619), the completeness (0.801) is 320% higher compared to manual labeling (0.190), and the F1 score (0.778) is 197% higher than the manual labeling (0.261).
((a))
((b))
((c))
@@
Generalizability. Our dataset consists of images collected from the lower extremities of human subjects. To qualitatively assess the generalizability of the trained model to new anatomies and specimens, we apply it to ultrasound images of the pelvis and femur collected from new specimens.Qualitative results of the trained model’s generalizability on new anatomies are shown in Figure 9
Figure 9: Examples of ultrasound images depicting the pelvis and femur, along with the corresponding predicted bone labels.
5 Discussion
The results highlight our methodology as a promising solution for collecting large-scale ultrasound datasets with automatically generated bone labels. Clinical evaluation results confirm the quality of these labels. Furthermore, the pelvis and hip experiment results demonstrate the effectiveness of the collected dataset in training deep learning models with strong generalizability.
Comparison with SOTA. Previous work [12] aimed to align a tracked CT bone model with tracked ultrasound images. However, they reported that the mismatch of 1–2 mm was too large to allow for accurate automatic bone labeling in ultrasound images, and manual labeling was adopted in the end. In contrast, the results show that our optimization pipeline enhances CT-ultrasound alignment, enabling automatic accurate bone label generation. In addition, unlike the traditional CT-US registration method [23], which calculates a single global transformation matrix for the entire ultrasound sweep to maximize overlapping voxel intensities, our approach computes a transformation matrix for each individual ultrasound frame. According to the results of the clinical evaluation, our method achieves significantly improved bone annotations. This improvement is statistically validated by the Wilcoxon signed-rank test with Bonferroni correction ($p<0.001$).
@@ 5 Discussion
Automatic vs Manual Labeling. Qualitative results in Figure 6 indicate that the model trained on UltraBones100k outperforms the surgeon in segmenting bones in ultrasound images. Manual labels mostly include only high-intensity bone regions. In contrast, predicted bone labels include both high-intensity and low-intensity regions. Quantitative results in Figure 8 demonstrate that the trained model consistently outperforms surgeon’s manual labeling, regardless of the distance threshold used. These qualitative and quantitative results in combination with the intensity distribution of labels illustrated in Figure 6(b) further confirm that the expert surgeon faces difficulty in correctly labeling bone contours in low-intensity regions. These findings underscore the superiority of our approach over the manual labeling.
Challenges in Dataset Availability. Unlike natural images, manually labeling bones in ultrasound images requires specialized expertise in ultrasound imaging and bone anatomy, making it financially and temporally expensive. This constraint is a key factor limiting the size of training datasets in existing studies [22], which typically range from only dozens to a few thousand samples. The small dataset size restricts the performance and generalizability of trained models. Moreover, most existing datasets are not publicly available. As highlighted in prior reviews [21, 22], the lack of a large-scale, publicly accessible dataset remains a major challenge in the ultrasound vision community. Our dataset collection methodology represents a solution to mitigate the dataset challenge. Furthermore, we make the proposed method, as well as the generated dataset publicly available. UltraBones100k is the largest known dataset with high quality bone labels, aiming to help advance future developments in intraoperative ultrasound for computer assisted surgery.
Generalizability. Although our dataset is the largest among existing works and includes specimens with diverse body conditions, its size remains relatively small compared to datasets in other domains. However, the fundamental principles of ultrasound imaging remain consistent across individuals, and the bone-soft tissue structure exhibits relatively uniform characteristics across different bodies. As a result, the appearance of bones in ultrasound images remains consistent. In the human body, bones are surrounded by soft tissues, and due to their high acoustic impedance, they consistently appear as bright regions in ultrasound images, regardless of anatomical differences. Additionally, since 2D ultrasound images capture only partial bone segments, they are often unaffected by variations in overall anatomical shape. The presence of shared structural features across bones—such as those in the arms and legs—further contributes to the model’s generalizability. Qualitative results in Figure 9 indicate the model’s generalizability on ultrasound images of the pelvis and femur, even though it was trained on data collected from the lower extremities.
Limitations. Although the experiment results demonstrate the generalizability of the model trained on UltraBones100k, its performance may decline in anatomies with significantly different bone-soft tissue structures and shapes, such as the spine and skull. Depending on the required level of accuracy, additional data collection may be necessary for specific anatomies and pathological conditions. In some cases, this may involve the collection of in vivo data. Conducting data collection studies with a limited number of patients is a common and justifiable approach in such scenarios. Furthermore, the publicly available model pretrained on UltraBones100k can be fine-tuned with newly collected data to enhance its performance for these specific cases.